Holland Brewery is considering adding a new piece of equipment that will improve labeling of bottles. There are two possibilities that are mutually exclusive with the following cash flows: Year Jiffy Labeler Super Labeler 0 -$17,100 -$22,430 1 5,100 7,500 2 5,100 7,500 3 5,100 7,500 4 5,100 7,500 5 5,100 7,500 Assume Holland Brewery’s WACC is 14 percent. Calculate the payback period, the NPV, the Profitability Index, the IRR, and the MIRR for both the Jiffy Labeler and the Super Labeler. Based on your calculations, which Labeler should be purchased?
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

L23 . А B | C / D | E | F | G | Year CF OHNM in Jiffy Labeler CF Cumulative CF ($17,100.00) ($17,100.00) $5,100.00 ($12,000.0

Cell reference -

DE F Jiffy Labeler Cumulative CF Year CE o =C4 -17100 5100 5100 5100 5100 5100 D4+C5 =D5+C6 =D6+C7 =D7+C8 =D8+c9 Super Labele

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.